Find the maximum and the minimum point of Y= -(X+1)^3 +8
1
Expert's answer
2019-09-20T11:05:20-0400

The derivative of "y" is "y'(x) = -3 (x+1)^2". Find critical points, by equating it to zero: "-3 (x+1)^2 = 0". The last equation has root "x = -1". The sign of the the derivative in both intervals "(-\infty, -1)" and "(-1, \infty)" is negative, so "x= -1" is not a relative extremum point. The range of function is "\mathbb{R}", so it doesn't have a maximum or minimum.